# Test-Time Adaptation with State-Space Models

### v1

@@ 3.3 Von Mises-Fisher Model for Hyperspherical Features
Choosing Gaussian densities for the transition and emission models, as discussed above, assumes the representation space follows an Euclidean geometry. However, prior work has shown that assuming the hidden representations lie on the unit hypersphere results in a better inductive bias for OOD generalization [26, 2]. This is due to the norms of the representations being biased by in-domain information such as class balance, making angular distances a more reliable signal of class membership in the presence of distribution shift [26, 2]. We too employ the hyperspherical assumption by normalizing the hidden representations such that $||{\mathbf{h}}||_{2}=1$ and model them with the von Mises-Fisher (vMF) distribution [25],
@@
where $\bm{\mu}_{k}\in{\mathbb{R}}^{D}$ with $||\bm{\mu}_{k}||_{2}=1$ denotes the mean direction of class $k$, $\kappa\in{\mathbb{R}}^{+}$ the concentration parameter, and $C_{D}(\kappa)$ the normalization constant. High values of $\kappa$ imply larger concentration around $\bm{\mu}_{k}$. The vMF distribution is proportional to a Gaussian distribution with isotropic variance and unit norm. While previous work [26, 27, 2] has mainly explored training objectives to encourage latent representations to be vMF-distributed, we apply Equation 7 to model the evolving representations.
@@
Returning to the SSM given above, we specify both transition and emission models as vMF distributions,
Hyperspherical transition model: $\quad p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1})=\prod_{k=1}^{K}{\text{{vMF}}}({\mathbf{w}}_{t,k}|{\mathbf{w}}_{t-1,k},\kappa^{\mathrm{trans}})$ (8)
Hyperspherical emission model: $\quad p({\mathbf{H}}_{t}|{\mathbf{W}}_{t})=\prod_{n=1}^{N_{t}}\sum_{k=1}^{K}\pi_{t,k}{\text{{vMF}}}({\mathbf{h}}_{t,n}|{\mathbf{w}}_{t,k},\kappa^{\mathrm{ems}})$ (9)
The parameter size of the vMF formulation only scales linearly with the feature dimension, i.e. $\mathcal{O}(DK)$ instead of $\mathcal{O}(D^{2}K)$ as for the Gaussian case. The noise parameters, $\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}}$ are scalar values, and we shall use the reduced parameter size to experiment with class conditioned noise parameters $\kappa_{k}^{\mathrm{trans}},\kappa_{k}^{\mathrm{ems}},k=1,\dots,K$ in Section 5. Figure 2 illustrates this STAD-vMF variant.
@@
Unlike in the linear Gaussian case, the vMF distribution is not closed under marginalization. Consequentially, the posterior distribution required for the expectation in Equation 5, $p({\mathbf{W}}_{1:T}{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$, cannot be obtained in closed form. We employ a variational EM objective, approximating the posterior with mean-field variational inference, following Gopal and Yang [16]:
@@
The variational distribution $q({\mathbf{W}},{\mathbf{C}})$ factorizes over $n,t,k$ and the objective from Equation 5 becomes $\operatorname*{arg\,max}_{\phi}\mathbb{E}_{q({\mathbf{W}},{\mathbf{C}})}\big{[}\log p({\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})\big{]}$. More details as well as the full maximisation steps for $\phi=\{\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}},\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K}\}$ can be found in Appendix B. Notably, posterior inference for this model is much more scalable than the Gaussian case, having operations that are linear in the dimensionality rather than cubic.
@@ 4 Related Work
Filtering for Deep Learning
Traditional filtering models, and the Kalman filter [19] in particular, have recently found use in deep learning as a principled way of updating a latent state with new information. In sequence modelling, filter-based architectures are used to learn the latent state of an observation trajectory in both discrete [23, 20, 13, 4] and continuous time [33, 1, 43]. However, here, the filtering model mimics the dynamics of individual observation sequences while we are interested in modeling the dynamics of the data stream as a whole. Chang et al. [7] and Titsias et al. [36] employ Kalman filters in a supervised online learning setting to update neural network weights to a non-stationary data stream. Like our method, Titsias et al. [36] infers the evolution of the linear classification head. However, [7, 36] rely on labels to update the latent state of the weights after prediction. In contrast, our weight adaptation is entirely label-free.
@@
Maintaining reliable predictions under distribution shift at test time has driven several research directions such as continual learning [10] and domain adaptation [30, 39]. Our setting falls into test-time adaptation, where the goal is to adapt the source-trained model given only access to the unlabeled target data [24, 41]. A simple yet effective approach is to keep updating the batch normalization (BN) statistics at test time [34, 28]. Based on this insight, a common strategy is to learn the parameters of the BN layer during test time [34, 28, 37, 15, 29, 31]. For instance, TENT [37] learns the BN parameters by minimizing entropy on the predicted target labels. A variation of the setting arises when the test distribution itself changes over time, a more realistic scenario studied by continual test-time adaptation. The main challenge in this paradigm is to ensure adaptability while preventing catastrophic forgetting of the source distribution. To mitigate this trade off, strategies involve episodic resetting to the source parameters [38, 31] or test sample selection [29]. Nonetheless, it has been observed that many strategies still experience performance degradation after extended periods of adaptation [29, 15, 38, 31].
@@ 5 Experiments
Our experimental goal is to demonstrate that STAD effectively models natural temporal drifts in the test-time covariates. This distinguishes our work from previous CTTA methods, which focus on domain shifts or shifts induced by corruption noise but do not consider natural evolution of time. Precisely, our setting of interest comprises (1) real-world shifts that (2) occur gradually over time and (3) lead to performance decay. Though ubiquitous in practice, systematic evaluation procedures for such drifts have only been considered fairly recently [40]. We use the evaluation protocol of the Wild-Time benchmark suite [40] to assess the adaptation performance of our model in this setting. In Section 5.1, we demonstrate that our method excels on natural temporal drifts yielding significant performance improvements in settings in which existing CTTA methods collapse. We show the importance of explicitly modeling shift dynamics via an ablation study in Section 5.2. Finally, in Section 5.3, we investigate limitations of our model and show how to diagnose failures ahead of time.
@@
We consider two image classification datasets exposed to natural temporal drifts. In addition to our main setting of interest, we also present results on a classic corruption dataset.
@@
  • Yearbook [14]: a dataset of portraits of American high school students taken across eight decades. Data shift in the students’ visual appearance is introduced by changing beauty standards, group norms, and demographic changes. We use the Wild-Time [40] pre-processing and evaluation procedure resulting into 33,431 images from 1930 to 2013. Each $32\times 32$ pixel, grey-scaled image is associated with the student’s gender as a binary target label. Images from 1930 to 1969 are used for training; the remaining years from 1970 to 2013 for testing.
@@
  • FMoW-Time: the functional map of the world (FMoW) dataset [22] maps 224 × 224 RGB satellite images to one of 62 land use categories. Distribution shift is introduced by technical advancement and economic growth changing how humans make use of land over time. FMoW-Time [40] is an adaptation from FMoW-WILDS [22, 8] that splits a total of 141,696 images into a training time period (2002 - 2012) and a testing time period (2013 - 2017).
@@
  • CIFAR-10-C: a dataset derived from CIFAR-10, to which 15 corruption / noise types are applied with 5 severity levels to introduce gradual distribution shift [17]. We increase the corruption severity starting from the lowest level (severity 1) to the most sever corruption (severity 5). This results in a test stream of $5\times 10,000$ images for each corruption type. Since our goal is mimicking gradual distribution shifts, we are not interested in switching between images of different corruption types, as previous work has done [38, 37].
Source Architectures
We use a variety of source architectures to demonstrate the model-agnostic nature of our method. They vary in parameter counts, backbone architecture and dimensionality of the representation space.
  • CNN: We employ the four-block convolutional neural network trained by [40] to perform the binary gender prediction on the yearbook dataset. Presented results are averages over three different random training seeds. The dimension of the latent representation space is 32.
  • DenseNet: For FMoW-Time, we follow the backbone choice of [22, 40] and use DenseNet121 [18] for the land use classification task. Weights for three random trainings seeds are provided by [40]. The latent representation dimension is 1024.
  • WideResNet: For the CIFAR-10 experiment, we follow [35, 37] and use the pre-trained WideResNet-28 [42] model from the RobustBench benchmark [9]. The latent representation have a dimension of 512.
Baselines
Despite the source model, we compare against three baselines suitable for an unsupervised, continuously-changing test stream. These baselines cover the most dominant paradigms in CTTA: adapting normalization statistics, entropy minimization and anti-collapse mechanics.
  • Source Model: the un-adapted original model.
  • BatchNorm (BN) Adaptation [34, 28]: aims to adapt the source model to distributions shift by collecting normalization statistics (mean and variance) of the test data.
  • Test Entropy Minimization (TENT) [37]: goes one step further and optimizes the BN transformation parameters (scale and shift) by minimizing entropy on test predictions.
  • Continual Test-Time Adaptation (CoTTA) [38]: takes a different approach by optimizing all model parameters with an entropy objective on augmentation averaged predictions and combines it with stochastic weight restore to prevent catastrophic forgetting.
5.1 Natural Temporal Distribution Drifts

### v2

@@ 3.3 Von Mises-Fisher Model for Hyperspherical Features
Choosing Gaussian densities for the transition and emission models, as discussed above, assumes the representation space follows an Euclidean geometry. However, prior work has shown that assuming the hidden representations lie on the unit hypersphere results in a better inductive bias for OOD generalization (Mettes et al., 2019; Bai et al., 2024). This is due to the norms of the representations being biased by in-domain information such as class balance, making angular distances a more reliable signal of class membership in the presence of distribution shift (Mettes et al., 2019; Bai et al., 2024). We too employ the hyperspherical assumption by normalizing the hidden representations such that $||{\mathbf{h}}||_{2}=1$ and model them with the von Mises-Fisher (vMF) distribution (Mardia & Jupp, 2009),
@@
where $\bm{\mu}_{k}\in{\mathbb{R}}^{D}$ with $||\bm{\mu}_{k}||_{2}=1$ denotes the mean direction of class $k$, $\kappa\in{\mathbb{R}}^{+}$ the concentration parameter, and $C_{D}(\kappa)$ the normalization constant. High values of $\kappa$ imply larger concentration around $\bm{\mu}_{k}$. The vMF distribution is proportional to a Gaussian distribution with isotropic variance and unit norm. While previous work (Mettes et al., 2019; Ming et al., 2023; Bai et al., 2024) has mainly explored training objectives to encourage latent representations to be vMF-distributed, we apply Eqn. 7 to model the evolving representations.
@@
Returning to the SSM given above, we specify both transition and emission models as vMF distributions, resulting in a hyperspherical transition model, $p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1})=\prod_{k=1}^{K}{\text{{vMF}}}({\mathbf{w}}_{t,k}|{\mathbf{w}}_{t-1,k},\kappa^{\mathrm{trans}})$, and hyperspherical emission model, $p({\mathbf{H}}_{t}|{\mathbf{W}}_{t})=\prod_{n=1}^{N_{t}}\sum_{k=1}^{K}\pi_{t,k}{\text{{vMF}}}({\mathbf{h}}_{t,n}|{\mathbf{w}}_{t,k},\kappa^{\mathrm{ems}})$. The parameter size of the vMF formulation only scales linearly with the feature dimension, i.e. $\mathcal{O}(DK)$ instead of $\mathcal{O}(D^{2}K)$ as for the Gaussian case. Notably, the noise parameters, $\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}}$ simplify to scalar values which reduces memory substantially. Fig. 3 illustrates this STAD-vMF variant.
@@
Unlike in the linear Gaussian case, the vMF distribution is not closed under marginalization. Consequentially, the posterior distribution required for the expectation in Eqn. 5, $p({\mathbf{W}}_{1:T}{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$, cannot be obtained in closed form. We employ a variational EM objective, approximating the posterior with mean-field variational inference, following Gopal & Yang (2014):
@@
The variational distribution $q({\mathbf{W}},{\mathbf{C}})$ factorizes over $n,t,k$ and the objective from Eqn. 5 becomes $\operatorname*{arg\,max}_{\phi}\mathbb{E}_{q({\mathbf{W}},{\mathbf{C}})}\big{[}\log p({\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})\big{]}$. More details including the full maximisation steps for $\phi=\{\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}},\{\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K}\}$ can be found in Sec. B.1. Notably, posterior inference for the vMF model is much more scalable than the Gaussian case. It operates with linear complexity in $D$, rather than cubic, reducing runtime significantly. Algorithm 1 (Sec. B.2) summarizes the method.
@@ 4 Related Work
State-Space Models for Deep Learning
Probabilistic state-space models, and the Kalman filter (Kalman, 1960) in particular, have found diverse use in deep learning as a principled way of updating a latent state with new information. In sequence modelling, filter-based architectures are used to learn the latent state of a trajectory in both discrete (Krishnan et al., 2015; Karl et al., 2017; Fraccaro et al., 2017; Becker et al., 2019) and continuous time (Schirmer et al., 2022; Ansari et al., 2023; Zhu et al., 2023). Inspired by traditional state-space models, recent work on structured state-space models (Gu et al., 2022; Smith et al., 2023; Gu & Dao, 2023) has pushed the state-of-the-art in sequence modeling. However, in this setting, the SSM models the dynamics for an individual sequence while we are interested in modeling the dynamics of the data stream as a whole. The latter is also the subject of supervised online learning, and Chang et al. (2023) and Titsias et al. (2023) employ Kalman filters in the supervised, non-stationary setting. Similarly to our method, Titsias et al. (2023) also infer with an SSM the evolution of the linear classification head. However, they rely on labels to update the latent state of the weights whereas our weight adaptation is entirely unsupervised.
@@
TTA improves model performance on unseen test data with distribution shifts by adapting to the test data during inference. Early TTA approaches focused on recalculating batch normalization (BN) statistics from test data (Nado et al., 2020; Schneider et al., 2020). This has often been combined with minimizing entropy (Liang et al., 2020; Wang et al., 2021; Zhang et al., 2022; Yu et al., 2024; Gao et al., 2024). Alternative objectives leverage contrastive learning (Chen et al., 2022), invariance regularization (Nguyen et al., 2023), and Hebbian learning (Tang et al., 2023). While many of these methods can be applied online, they don’t fully address the challenges of non-stationary test streams. CTTA aims to handle changing distributions without forgetting source knowledge, using strategies such as episodic resets (Press et al., 2024), student-teacher models (Wang et al., 2022; Döbler et al., 2023; Brahma & Rai, 2023), masking (Liu et al., 2024), and regularization (Niu et al., 2022; Song et al., 2023). However, temporally correlated data remains challenging as they can cause online class imbalance and disrupt BN statistics, leading to model collapse. Solutions include adapted BN strategies (Zhao et al., 2023a; Lim et al., 2023), reservoir sampling (Gong et al., 2022; Yuan et al., 2023), filtering unreliable samples (Niu et al., 2023), and tracking label distributions (Zhou et al., 2023). Methods that bypass BN and instead adapt the classification head avoid collapse under temporal correlation. These methods adjust logits (Boudiaf et al., 2022) and predict based on KNNs (Jang et al., 2023). Most similar to our method, T3A (Iwasawa & Matsuo, 2021) recomputes prototypes from representations. However, T3A relies upon heuristics while STAD explicitly models dynamics using a SSM.
@@ 5 Experiments
We evaluate our method, STAD, against various baselines on a range of datasets under challenging settings. In Sec. 5.1, we study temporal distribution shifts as defined in Sec. 2, demonstrating the difficulty of the task and STAD’s robustness in practical settings. In Sec. 5.2, we go beyond temporal shifts and find that STAD is competitive on reproduction datasets and synthetic corruptions as well. Finally, in Sec. 5.3, we provide insights into STAD’s mechanisms, confirming the reliability of its prototypes and highlighting the importance of modeling shift dynamics through an ablation study. We now describe the datasets, source architectures, and baselines. Further implementation details can be found in App. C.
@@
Our primary focus is on temporal distribution shifts. To further evaluate the effectiveness of our method, we also test its performance on reproduction datasets and standard image corruptions.
@@
  • Yearbook (Ginosar et al., 2015): a dataset of portraits of American high school students taken across eight decades. Data shift in the students’ visual appearance is introduced by changing beauty standards, group norms, and demographic changes. We use the Wild-Time (Yao et al., 2022) pre-processing and evaluation procedure resulting into 33,431 images from 1930 to 2013. Each $32\times 32$ pixel, grey-scaled image is associated with the student’s gender as a binary target label. Images from 1930 to 1969 are used for training; the years 1970 - 2013 for testing.
  • EVIS: the evolving image search (EVIS) dataset (Zhou et al., 2022) consists of images of 10 electronic product and vehicle categories retrieved from Google search, indexed by upload date. The dataset captures shift caused by rapid technological advancements, leading to evolving designs across time. It includes 57,600 RGB images of 256x256 pixels from 2009 to 2020. Models are trained on images from 2009-2011 and evaluated on images from 2012-2020.
@@
  • FMoW-Time: the functional map of the world (FMoW) dataset (Koh et al., 2021) maps $224\times 224$ RGB satellite images to one of 62 land-use categories. Distribution shift is introduced by technical advancement and economic growth changing how humans make use of the land. FMoW-Time (Yao et al., 2022) is an adaptation of FMoW-WILDS (Koh et al., 2021; Christie et al., 2018), splitting 141,696 images into a training period (2002-2012) and a testing period (2013-2017).
  • CIFAR-10.1 (Recht et al., 2019): a reproduction of CIFAR-10 (Krizhevsky et al., 2009) assembled from the same data source by following the same cleaning procedure. The dataset contains 2,000 $32\times 32$ pixel images of 10 classes. Models are trained on the original CIFAR-10 train set.
  • ImageNetV2 (Recht et al., 2019): a reproduction of ImageNet (Deng et al., 2009) with 10,000 images of 1,000 classes scaled to $224\times 224$ pixels. Models are trained on the original ImageNet.
@@
  • CIFAR-10-C: a dataset derived from CIFAR-10, to which 15 corruption types are applied with 5 severity levels (Hendrycks & Dietterich, 2019). We mimic a gradual distribution shift by increasing the corruption severity starting from the lowest level (severity 1) to the most sever corruption (severity 5). This results in a test stream of $5\times 10,000$ images per corruption type.
Source Architectures and Baselines
We employ a variety of source architectures to demonstrate the model-agnostic nature of our method. They notably vary in backbone architecture (we use CNN, DenseNet, ResNet, WideResNet) and dimensionality of the representation space (from 32 up to 2048). We list details in Sec. C.1. In addition to the source model, we compare against 7 baselines representing fundamental approaches to TTA. Five of them adapt the feature extractor: Batch norm adaptation (BN) (Schneider et al., 2020; Nado et al., 2020), TENT (Wang et al., 2021), CoTTA (Wang et al., 2022), SHOT (Liang et al., 2020) and SAR (Niu et al., 2023). Like our method STAD, two baselines adapt the last linear layer: T3A (Iwasawa & Matsuo, 2021) and LAME (Boudiaf et al., 2022). More details are provided in Sec. C.2. Batch sizes are the same for all baselines (Sec. C.3). To ensure optimal performance on newly studied datasets, we conduct an extensive hyperparameter search for each baseline, following the hyperparameters and value ranges suggested in the original papers. We then report the best settings. Sec. C.3 specifies the grid searches we conducted.
5.1 Temporal Distribution Shifts
We start by evaluating the adaptation abilities to temporal distribution shift on three image classification datasets (Yearbook, EVIS, FMoW-Time), which vary in number of classes (2, 10, 62, respectively), representation dimension (32, 512, 1024, respectively) and shift dynamics (recurring, progressive and rapid, respectively as visible in Fig. 4). For the low-dimensional representations of Yearbook, we also evaluate our computationally costly Gaussian model (STAD-Gauss). We evaluate two settings: (i) covariate shift with a uniform label distribution and (ii) covariate shift with additional shift in the label distribution $\mathbb{Q}_{t}({\mathbf{y}})$. Having a uniform label distribution—samples are evenly shuffled, making test batches nearly class balanced—has been the standard evaluation setting for TTA. However, particularly in temporal distribution shifts, it is highly unlikely that samples arrive in this iid-manner. Instead, temporally correlated test streams often observe consecutive samples from the same class (Gong et al., 2022). We follow Lim et al. (2023), ordering the samples by class and thus inducing an extreme label shift. We draw the class order uniformly at random.
Table 2: Accuracy on temporal distribution shifts and label shifts, averaged over three random seeds. Colors highlight performance gains and drops relative to the source model. Best model in bold, second-best underlined.
Yearbook EVIS FMoW-Time
Method covariate shift + label shift covariate shift + label shift covariate shift + label shift
Source model 81.30 $\pm$ 4.18 56.59 $\pm$ 0.92 68.94 $\pm$ 0.20
adapt feature extractor
BN 84.54 $\pm$ 2.10 70.47 $\pm$ 0.33 45.72 $\pm$ 2.79 14.48 $\pm$ 1.02 67.60 $\pm$ 0.44 10.14 $\pm$ 0.04
TENT 84.53 $\pm$ 2.11 70.47 $\pm$ 0.33 45.73 $\pm$ 2.78 14.49 $\pm$ 1.02 67.86 $\pm$ 0.54 10.21 $\pm$ 0.01
CoTTA 84.35 $\pm$ 2.13 66.12 $\pm$ 0.87 46.13 $\pm$ 2.86 14.71 $\pm$ 1.00 68.50 $\pm$ 0.25 10.19 $\pm$ 0.04
SHOT 85.17 $\pm$ 1.89 70.71 $\pm$ 0.20 45.93 $\pm$ 2.75 14.51 $\pm$ 1.00 68.02 $\pm$ 0.51 10.08 $\pm$ 0.07
SAR 84.54 $\pm$ 2.10 70.47 $\pm$ 0.33 45.78 $\pm$ 2.80 14.63 $\pm$ 1.00 67.87 $\pm$ 0.51 10.27 $\pm$ 0.10
adapt classifier
LAME 81.60 $\pm$ 3.99 82.70 $\pm$ 4.55 56.67 $\pm$ 0.99 69.37 $\pm$ 5.37 68.32 $\pm$ 0.32 83.05 $\pm$ 0.48
T3A 83.49 $\pm$ 2.55 83.46 $\pm$ 2.59 57.63 $\pm$ 0.77 57.32 $\pm$ 0.77 66.77 $\pm$ 0.26 66.83 $\pm$ 0.27
STAD-vMF (ours) 85.50 $\pm$ 1.34 84.46 $\pm$ 1.19 56.67 $\pm$ 0.82 62.08 $\pm$ 1.11 68.87 $\pm$ 0.06 86.25 $\pm$ 1.18
STAD-Gauss (ours) 86.22 $\pm$ 0.84 84.67 $\pm$ 1.46
Figure 4: Accuracy over time for three temporal distribution shifts: On Yearbook for example, STAD can mitigate the distribution shift by improving 10 points upon the source model for some years (1980s). Some baselines perform similarly, which is depicted by overlaying accuracy trajectories. Error bars for FMoW-Time are tiny.
Temporal shifts pose challenges for existing TTA methods.
Tab. 2 shows overall accuracy, averaged over all time steps and three random training seeds. Results that do not outperform the source model are highlighted in red and ones that do in blue. Methods that primarily adapt the feature extractor are shown in the upper section of the table. Ones that, like ours, adapt the classifier are shown in the lower section. To summarize the results: on Yearbook, all methods perform well without label shift, and with label shift, only classifier-based methods improve upon the source baseline. Feature-based methods completely fail on EVIS, and all models, except LAME and STAD-vMF under label shift, fail on FMoW-Time. This leads us to three key takeaways: first, these TempTTA tasks are inherently difficult, leading to smaller adaptation gains overall compared to traditional corruption experiments. Second, methods that adapt only the last layer clearly perform better on temporal distribution shifts under both label distribution settings. This indicates that perhaps ‘less is more’ for TempTTA. Third, STAD demonstrates the most consistent performance, ranking as the best or second-best model across all datasets and settings. On Yearbook, both the Gaussian and vMF variants outperform the baselines, with the fully parameterized Gaussian model better capturing the distribution shift than the more lightweight vMF model. Fig. 4 displays adaptation performance over different timestamps. We see that on EVIS (middle) the methods markedly separate, which reflects the aforementioned gap between feature-based and classifier-based approaches. The reader may also wonder if STAD can be stacked on top of a feature-based approach. We present results exemplary for BN in Tab. 8 (Sec. D.2) but found that it offers no significant improvement to STAD’s performance.
STAD excels under label shift
Tab. 2 demonstrates that STAD performs particularly well under imbalanced label distributions, delivering the best results on both Yearbook and FMoW-Time. This advantage stems from STAD’s clustering approach, where a higher number of samples from the same ground truth class provides a stronger learning signal, leading to more accurate prototype estimates. This is particularly notable on FMoW, where STAD improves upon the source model by more than 17 points. Further, the performance gap between classifier and feature extractor adaptation methods becomes even more pronounced in this setting. This is not surprising as the latter typically depend heavily on current test-batch statistics, making them vulnerable to imbalanced class distributions (Niu et al., 2023). In contrast, having fewer classes to cause confusion allows classifier-based methods to benefit from label shift, with STAD delivering the most persistent adaptation gains.
Figure 5: Batch size effects under covariate shift (first row) and additional label shift (second row): Compared to most baselines, STAD-vMF (dark blue) is relatively robust to very small batch sizes. For label shift on EVIS and FMoW-Time, we observe a sweet spot around batch size 16.
STAD is robust to small batch sizes
Adapting to a small number of samples is crucially valuable, as one does not have to wait for a large batch to accumulate in order to make adapted predictions. We next evaluate performance across 12 different batch sizes ranging from 1 to 2048 under both covariate shift and additional label shift. Fig. 5, displays results. STAD-vMF (dark blue line) is able to maintain stable performance under all batch sizes. In Tab. 7 (Sec. D.1), we report values for batch size 1 showing that STAD adapts successfully even in the most difficult setting. In contrast, methods relying on normalization statistics collapse when not seeing enough samples per adaptation step. For example, on FMoW, feature-based methods collapse to nearly random guessing at the smallest batch sizes. When batch sizes are large, note that TENT, CoTTA, SHOT, and SAR hit memory constraints on FMoW-Time, failing to close the gap to the source model performance in the label shift scenario.
5.2 Beyond Temporal Shifts: Reproduction Datasets and Synthetic Corruptions
Although STAD is designed for temporal distribution shifts, we are also interested in the applicability of our method to other types of shifts. Next we report performance on reproduction datasets and synthetic image corruptions.